# Forbidden Facts: An Investigation of Competing Objectives in Llama-2

### 14 Dec 2023 v1

@@ 1 Forbidden fact task
In this paper, we study the forbidden fact task, where a model is instructed to truthfully perform factual recall under the constraint of not being able to say a forbidden word. We use the following prompt template (adapted for Llama-2-chat models [13]):
@@ 2 Decomposing Llama-2
$\mathrm{next\_token\_distribution}(\mathbf{p})=\mathrm{Softmax}\left(W_{U}\cdot\mathrm{LayerNorm}\left(\ \sum_{i}r_{i}(\mathbf{s})\right)\right),$ (2.1)
@@
We will attribute the suppression behavior of Llama-2 on competing prompts to these residual stream components, of which there are 1057 in total: 1 initial embedding component, 1024 attention head components (32 heads per layer over 32 layers), and 32 MLP head components (one per layer). This decomposition is quite natural with respect to the standard way transformers are implemented.
@@ 3 Analysis of most important components
Out of the 35 components that comprise the aggregregate circuit, 31 are attention heads and 10 are MLPs. We prioritized attention head analysis since we seek to understand the information flow of the model across tokens. In future work, we will attempt to decompose the inputs to the MLPs and use sparse-probing and autoencoders to interpret features from individual neurons [5, 1].
@@
Figure 3.2: From left to right, examining L18H9: 1) Fixing the final query vector, we plot the attention to the forbidden tokens (which can be correct, relevant incorrect, and irrelevant incorrect) as they are enriched through the layers of the model. Note that the key enrichment is semantically specific. 2) Fixing the forbidden key vectors, we plot the attention to the forbidden tokens as the query is enriched through the layers of the model. 3) We patch the query token for the competing run with a corresponding uniformly sampled noncompeting query: the x-axis represents the log odds of the attention to the forbidden token for the competing run, and the y-axis represents the same quantity for the noncompeting runs. Note that these values are strongly correlated, so the query enrichment is not semantically specific. 4) Here we shuffle the positional embeddings instead of patching to determine the y-axis. Note that these values are also strongly correlated, so query enrichment is not positionally specific. See Figure G.1 for data on more heads.
We define enrichment as the process by which a particular embedding gains information as it moves through the model layers. Broadly, we find that both the key and query enrichment help the suppressor heads attend to the forbidden token for the competing prompt (3.2). This motivates the questions of how this information arrives to the key and query vectors.
Interestingly, we find that key attention is not positionally specific, and query attention is not semantically specific. To test invariance to position, we randomly shuffle the positional embeddings of the keys. To test invariance to semantic content, we patch the query vectors from competing runs with the corresponding query vectors from random noncompeting queries.
We find, surprisingly, that the only type of specificity the suppressor heads exhibit is key semantic specificity: the heads privilege correct answers to the factual recall over all other keys. This is consistent with the hypothesis that one of the primary mechanisms the model uses to attend to the correct key is the "knowledge" that the key token is the correct completion to the fact.
Because of this, we believe that the model uses a more complex mechanism to communicate what to suppress to the suppressor heads than simply directly suppressing the tokens specified by the forbidden token instruction. Furthermore, Figure F.2 and Figure G.1 show that there exists significant heterogeneity in attention enrichment behavior as well.
@@ 4 Discussion
If even simple behaviors for which there exist trivial algorithmic solutions can be implemented by AI systems in complex, heterogeneous ways, then achieving the first two goals may be extremely difficult advanced AI systems. This is a sobering state of affairs. However, we discuss two possible ways that this conclusion need not follow from our results.
@@
Because query enrichment is not specific to the forbidden word (Section 3), we find words these heads preferentially attend to. For example, L27H29 prefers paying attention to "California". On a noncompeting run for the factual recall task "The Golden Gate Bridge is in the state of" with an irrelevant forbidden word "floor", the model responds correctly with "California" 96.3% of the time. However, by adding two words to the first sentence of the prompt: "You are an obedient assistant from California who only responds with a single word with no punctuation.", we can completely break the model.
In particular, this combination of forbidding an irrelevant word and placing "California" innocuously in the system prompt leads the suppressor components to pay attention and suppress "California" to a 17.7% completion probability, elevating "San Francisco", an incorrect answer, to the top response. We also find that patching just the top suppressor head from the noncompeting run results in "California" again being the top answer, at 37.3%.
@@ Acknowledgements
Tony Wang and Miles Wang were supported by the Existential Risk Laboratory, and Kaivu Hariharan was supported by the MIT Tang Family FinTech Undergraduate Research and Innovation Scholarship as part of the MIT SuperUROP program.
@@ Appendix D Why log-odds and log Bayes factors?
So the log-odds of 99% is $\log_{10}(.99/(1-.99))=\log_{10}99\approx 2$ dits and the log-odds of 99.9999% is $\log_{10}(.999999/(1-.999999)=\log_{10}999999\approx 6$ dits. Thus in order to transform 99% to 99.9999%, we have apply an update with a log Bayes factor of approximately 4 dits. Likewise, if one does the math, the log Bayes factor needed to change 1% to 0.00001% is precisely the negative of the above, or roughly $-4$ dits.
@@ Appendix G Key query specificity plots
Figure G.1: More data for Llama-2-7b-chat in the format of Figure 3.2.
@@ Appendix H California Attacks
Because query enrichment is not specific to the forbidden word, we find words these heads preferentially attend to. For example, L27H29 prefers paying attention to "California". On a noncompeting run for the factual recall task "The Golden Gate Bridge is in the state of" with irrelevant forbidden word "floor", the model responds correctly with "California" at 96.3%. A California Attack simply adds two words to the first sentence of the prompt: "You are an obedient assistant from California who only responds with a single word with no punctuation."

### 31 Dec 2023 v3

@@ 1 Forbidden fact task
In this paper, we study the forbidden fact task, where a model is instructed to truthfully perform factual recall under the constraint of not being able to say a forbidden word. We use the following prompt template (adapted for Llama-2-chat models [^1] Results are for Llama-2-7b-chat unless stated otherwise. See the Appendix for results on other models sizes. [13]):
@@ 2 Decomposing Llama-2
$\mathrm{next\_token\_distribution}(\mathbf{p})=\mathrm{Softmax}\left(W_{U}\cdot\mathrm{LayerNorm}\left(\ \sum_{i}r_{i}(\mathbf{p})\right)\right),$ (2.1)
@@
We will attribute the suppression behavior of Llama-2 on competing prompts to these residual stream components. For the 7b model, there are 1057 components in total: 1 initial embedding component, 1024 attention head components (32 heads per layer over 32 layers), and 32 MLP head components (one per layer). This decomposition is quite natural with respect to the standard way transformers are implemented.
@@ 3 Analysis of most important components
Out of the 35 components that comprise the aggregregate circuit, 28 are attention heads and 7 are MLPs. We prioritized attention head analysis since we seek to understand the information flow of the model across tokens. In future work, we will attempt to decompose the inputs to the MLPs and use sparse-probing and autoencoders to interpret features from individual neurons [5, 1].
@@
Figure 3.2: Behavior of attention head L18H9 over the Forbidden Facts dataset (lines denote median quantities). From left to right: 1) We fix the query-vector of L18H9 at the final token position and plot how much it attends to partially enriched key-vectors of the forbidden word tokens. Partially enriched key-vectors are generated by feeding output activations of early layers into the key-circuit of L18H9. Key-enrichment by earlier layers is critical for achieving the full attention effect. 2) We fix the key-vectors of the forbidden word tokens, and plot the attention paid to them by partially enriched final-token query-vectors. Partial enrichment is done in the same manner as in the first plot. Query-enrichment is also critical for achieving the full attention effect. 3) We analyze whether the final-token query-vector from a competing run (where the correct answer is forbidden) will attend to key-vectors from a non-competing run (where a word other than the correct-answer is forbidden). We compare this against the baseline of how much the final token attends to the forbidden word in unmodified competing runs. The x-axis shows the log-odds of attention in the baseline run, and the y-axis shows the log-odds of attention in our cross-run experiment. The strong correlation indicates the attention mechanism is not semantically specific. 4) The x-axis is the same as in plot 3, and the y-axis is attention paid by the final-token query-vector to forbidden-token key-vectors when we randomize the position-embedding of the input activations to the key-circuit. The strong correlation indicates the attention mechanism is not positionally specific. See Figure G.1 for data on more heads.
We define enrichment as the process by which a particular embedding gains information as it moves through the model’s layers [4]. In Figure 3.2, we find that both key and query enrichment help the suppressor heads attend to the forbidden token. We also find that the attention paid by suppressor heads to the forbidden word is neither positionally nor semantically specific. Positional non-specificity means that the attention pattern does not depend on the position encoding of the forbidden-word activations. Semantic non-specificity means the attention pattern does not depend on what we set the forbidden word as.
Overall, despite a moderate amount of effort, we were unable to fully reverse-engineer the exact attention mechanism for even a single suppressive attention head. Furthermore, Figure F.2 and Figure G.1 show that there exists significant heterogeneity in attention enrichment behavior as well. In the next section, we give some evidence (in the form of the California Attack) that our observations may be due to the model’s mechanisms being inherently complicated.
@@ 4 Discussion
If even simple behaviors for which there exist trivial algorithmic solutions can be implemented by AI systems in complex, heterogeneous ways, then achieving the first two goals may be extremely difficult for advanced AI systems. This is a sobering state of affairs. However, we discuss two possible ways that this conclusion need not follow from our results.
@@
In our analysis of suppressive attention heads, we found that these heads could be tricked into attending to words that were not the forbidden word. For example, attention head L27H29 likes paying attention to the "California" token even when it is not the forbidden word. On a noncompeting run for the factual recall task "The Golden Gate Bridge is in the state of" with an irrelevant forbidden word "floor", the model responds correctly with "California" 96.3% of the time. However, by adding two words to the first sentence of the prompt: "You are an obedient assistant from California who only responds with a single word with no punctuation.", we can break the model.
In particular, this combination of forbidding an irrelevant word and placing "California" innocuously in the system prompt leads the suppressor components to pay attention and suppress "California" to a 17.7% completion probability, elevating "San Francisco", an incorrect answer, to the top response. We also find that patching just the top suppressor head from the noncompeting run results in "California" again being the top answer, at 37.3%. We found that California Attacks worked for both Llama-2-7b-chat and Llama-2-13b-chat. See Appendix H for more details.
@@ Acknowledgements
Tony Wang and Miles Wang were supported by the Existential Risk Laboratory, and Kaivalya Hariharan was supported by the MIT Tang Family FinTech Undergraduate Research and Innovation Scholarship as part of the MIT SuperUROP program.
@@ Appendix D Why log-odds and log Bayes factors?
So the log-odds of 99% is $\log_{10}(.99/(1-.99))=\log_{10}99\approx 2$ dits and the log-odds of 99.9999% is $\log_{10}(.999999/(1-.999999)=\log_{10}999999\approx 6$ dits. Thus in order to transform 99% to 99.9999%, we have to apply an update with a log Bayes factor of approximately 4 dits. Likewise, if one does the math, the log Bayes factor needed to change 1% to 0.00001% is precisely the negative of the above, or roughly $-4$ dits.
@@ Appendix G Key query specificity plots
Figure G.1: More data for Llama-2-7b-chat in the format of Figure 3.2. See Figure G.2 for the Llama-2-13b-chat version of this plot.
@@ Appendix H California Attacks
Because suppressor heads can pay attention to words that are not the forbidden word, we construct adversarial examples by finding words these heads preferentially attend to. For example, L27H29 prefers paying attention to "California". On a noncompeting run for the factual recall task "The Golden Gate Bridge is in the state of" with irrelevant forbidden word "floor", the model responds correctly with "California" at 96.3%. A California Attack simply adds two words to the first sentence of the prompt: "You are an obedient assistant from California who only responds with a single word with no punctuation."